# A Case for Library-Level k-Means Binning in Histogram Gradient-Boosted Trees

### v1

@@ Abstract
Modern gradient-boosted decision trees (GBDTs) accelerate split finding with histogram-based binning, which reduces complexity from O(N) to O(B) given a fixed bin budget B. However, the predominant quantile binning strategy—designed to distribute data points evenly among bins—may overlook critical boundary values that could enhance predictive performance. In this work, we propose replacing quantile binning with a k-means discretizer initialized with quantile bins. We test this swap on 33 OpenML tasks plus synthetics that control for modality, skew, and bin budget. Across 18 regression datasets, k-means shows no statistically significant losses at the 5% level and wins in four cases—most strikingly a 55% MSE drop on one particularly skewed dataset—even though k-means’ mean reciprocal rank (MRR) is slightly lower (0.65 vs 0.72). On the 15 classification datasets the two methods are statistically tied (MRR 0.70 vs 0.68) with gaps 0.2 pp. Synthetic experiments confirm consistently large MSE gains—typically >20% and rising to 90% as outlier magnitude increases or bin budget drops. We find that k-means keeps error on par with exact splitting when extra cuts add little value, yet still recovers key split points that quantile overlooks. As such, we advocate for a built-in bin_method=k-means flag, especially in regression tasks and in tight-budget settings such as the 32–64-bin GPU regime—because it is a "safe default" with large upside, yet adds only a one-off, cacheable overhead ( 2s to bin 10M rows on one core).
@@ 1 Introduction
Gradient Boosted Decision Trees (GBDTs) are ensemble learning methods that have achieved state-of-the-art results on a wide range of tasks. By iteratively fitting new decision trees to the "pseudo-residuals" of a loss function, GBDTs combine many weak learners (e.g. individual decision trees) into a strong predictor (Friedman, 2001). Since the seminal work of Friedman on gradient boosting, the technique has become a de facto choice for structured data problems. Because GBDTs require only a twice-differentiable loss function, they are especially useful for non-standard prediction problems, such as ranking or quantile regression (Burges, 2010). Modern implementations like XGBoost, LightGBM, and CatBoost have popularized GBDTs by offering order-of-magnitude computational speed-ups with excellent accuracy (Chen and Guestrin, 2016) (Ke et al., 2017) (Prokhorenkova et al., 2019). These systems have been widely adopted to win Kaggle and other ML competitions, often outperforming deep learning on tabular data (Grinsztajn et al., 2022).
@@
A key method for optimizing these algorithms is their use of histogram-based binning for splitting continuous features. Traditional decision tree algorithms sort feature values to find split points, which is computationally expensive ( per feature initially, and afterwards) and thus intractable for large datasets. To allow training on datasets sometimes too large to fit into memory, modern GBDT algorithms bucket continuous feature values into discrete bins and accumulate gradients per bin during training. This approximation drastically reduces computation and memory overhead while maintaining similar accuracy as exact splitting. While initial histogram-binning does require one scan over all features, split-finding only requires O(#bins) per feature. When #bins #data, histogram-based algorithms can achieve order-of-magnitude speedups. Major state-of-the-art algorithms, like XGBoost, LightGBM, and CatBoost, all use similar histogram-binning methods, though with small differences in implementation. These innovations allow GBDTs to effectively scale to datasets with millions or billions of instances.
@@
In state-of-the-art histogram-based GBDT libraries, the default choice is quantile binning: thresholds are placed so that every bin holds roughly the same number of observations, barring duplicate values. This equal-frequency distribution is popular because it preserves the rank order of values and often yields split quality comparable to those obtained by exact algorithms. However, quantile binning can sometimes mask rare but influential values. For example, in a sample of one million observations with one thousand extreme outliers, a 255-bin quantile scheme would pool those outliers with 3000 ordinary points, blunting a potentially important boundary.
@@
A natural alternative is to utilize some method of unsupervised clustering to isolate important points prior to training. One such example is via 1-D k-means clustering (Lloyd, 1982), which explicitly minimizes within-bin variance and can place cuts at rare yet influential values. We therefore compare k-means clustering with quantile initialization111 Although Wang and Song (2011) proved there exist 1-D k-means clustering algorithms that guarantee optimality, we find the computational requirements are too heavy and thus choose to utilize the well-known iterative Lloyd k-means algorithm. with quantile binning as well as uniform (equal-width) binning to characterize the regimes—tail mass, multi-modality, and bin budget—under which each discretizer is preferable and to test whether a simple k-means swap can boost accuracy without compromising training speed.
@@
We find that uniform equal-width cuts seldom rival quantile, whereas k-means almost always draws level with quantile and occasionally provides a marked boost. Across 18 real-world regression datasets, k-means never performs worse than quantile at the 5% level and cuts MSE by more than half on the most skewed case, though it has a slightly lower MRR (0.65 vs 0.72). When we repeat the sweep with the 63-bin budget recommended for GPU training (Zhang et al., 2017), k-means posts four wins and raises its best-case drop to 68% MSE (Table 4), though quantile achieves two small () wins. In 15 classification datasets quantile and k-means are statistically tied (MRR 0.68 vs 0.7) with gaps pp–differences that are essentially negligible. Synthetic regression benchmarks tell the same story: when outliers dominate or when bin budget is low, k-means often cuts error by more than 20% and can reduce error by as much as 90% in some cases. Taken together, our findings show that k-means stays within statistical noise of exact splitting when the bin budget is generous, yet often uncovers critical split points that quantile misses.
The extra cost—about 0.3s to bin one million rows on a single core, or 2.5s to bin ten million rows—is a one-off, cacheable operation.
To our knowledge this is the first systematic comparison of binning schemes for GBDTs, and the results highlight a simple, practical way for commercial GBDT libraries to tighten performance without altering downstream training code.
@@ 2 Related Work
In recent years, gradient-boosted decision trees (GBDTs) have become the dominant choice for large-scale tabular prediction thanks to highly optimized "histogram" implementations that reduce split search from O(n) to O(# bins) per feature by pre-binning continuous inputs. LightGBM introduced leaf-wise growth together with GPU-friendly, histogram-based training and a gradient-based one-side sampling scheme to speed up the scan over candidate bins (Ke et al., 2017). XGBoost introduces a weighted quantile sketch to obtain approximate split points that are -accurate even on distributed data (Chen and Guestrin, 2016). CatBoost further refines this pipeline with ordered boosting and efficient handling of categorical features, yielding state-of-the-art accuracy on many public benchmarks (Prokhorenkova et al., 2019). Scikit-learn, a package for easy application of ML algorithms, recently added HistGradientBoosting as a re-implementation of these histogram algorithms for GBDTs, providing a convenient baseline for academic comparisons (Pedregosa et al., 2018).
Although the histogram abstraction is now standard, the choice of binning strategy has received surprisingly little dedicated study. All major libraries employ equal-frequency (quantile) binning; the only divergences are LightGBM’s gradient-based one-side-sampling and XGBoost’s quantile sketch that prioritize high-loss regions (Ke et al., 2017)(Chen and Guestrin, 2016).
Adaptive or data-driven binning has been explored in other settings. Early entropy-based cuts (Fayyad and Irani, 1993) and MDL-guided binning (Dougherty et al., 1995) found that supervised partitions improve decision-tree and Naïve-Bayes accuracy over equal-width binning. In federated settings, Ong et al. (2020) begin with equal-mass bins and dynamically merge or split them using gradient entropy to keep communication costs fixed. Our work is orthogonal to these settings: we study library-level, label-agnostic binning that can serve as the starting point for any histogram-GBDT pipeline, federated or centralized. Although k-means is unsupervised, its variance-minimizing objective lets it capture label-relevant tails that equal-mass schemes often smooth over, thus recovering a portion of the gains previously attributed only to supervised or adaptive methods.
@@ 3 Experiments
We evaluate our discretizers on 33 OpenML tasks and a suite of controlled synthetics. The complete reproducibility package—source code, raw logs, and result tables—is openly available at github.com/ashlab11/KMeansGBDT under the Apache 2.0 licence.
@@ 3.1.1 Methodology
To ensure replicable results we evaluate our models on the OpenML (Vanschoren et al., 2014) benchmark suite described in Grinsztajn et al. (2022) (study_id 336 for regression, 337 for classification), dropping one task from each track (Higgs, Zurich Delays) due to computational constraints. The remaining 18 regression and 15 classification tasks span – instances and 2–420 numeric features. Appendix 2 lists observations and features for each dataset.
@@
We compare three discretizers: quantile (LightGBM default), uniform (equal-width), and -means (Lloyd with quantile seeding). Unless noted otherwise, all binning schemes use bins (the common default in LightGBM and scikit-learn, among others).222Appendix 4 reports an ablation at the GPU-recommended budget of , as described in Zhang et al. (2017). We also compare these discretizers with a naive approach that checks all possible split points.
@@
Commercial-grade libraries such as LightGBM, XGBoost, and CatBoost apply additional preprocessing layers—e.g. LightGBM’s exclusive-feature bundling, or XGBoost’s weighted quantile sketch (Ke et al., 2017) (Chen and Guestrin, 2016). These steps are re-executed even when the input has already been pre-binned to 255 distinct values, producing a second, library-specific histogram that blurs the effect we wish to measure. To isolate the contribution of the external discretizer itself, we therefore run all main-paper experiments with scikit-learn’s vanilla GradientBoostingRegressor/Classifier, where the model consumes our bins exactly as supplied. In a supplementary run with XGBoost (Appendix 5)—configured to bypass its internal quantile sketch—we observed similar results on representative regression datasets, suggesting that the choice of baseline model does not influence our conclusions.
For every (dataset, binning, learner) triple we run a 30-trial RandomizedSearchCV with 5-fold CV on hyper-parameters shown in Appendix 3. Each experiment is repeated over 20 random train/test splits (80/20) to estimate variability.
@@
We report mean squared error (MSE) for regression and ROC-AUC for classification. Regression MSE values span several orders of magnitude, so each row of our results table presents them in scientific notation with the common exponent factored out on the left of the dataset name (e.g. "Brazilian Houses "). This keeps the numeric columns directly comparable across datasets.
For each dataset, we compute a custom mean-reciprocal-rank (MRR): for each dataset the three histogram methods (the exact method is excluded) are ranked; the reciprocal rank is then averaged over datasets.
Paired two-sided t-tests (n=20 splits) compare the top-ranked discretizer with the runner-up (exact excluded) on each dataset. All resulting p-values are jointly adjusted with the Benjamini–Hochberg procedure (Benjamini and Hochberg, 1995); cells that stay significant at the level after this correction are marked **.
@@ 3.1.2 Results
Tables 2(a) and 2(b) report mean test-set scores averaged over 20 random 80:20 splits, with per-dataset winners flagged by paired t-tests, as described in section 3.1.1.
Regression (18 tasks) k-means attains the best mean-reciprocal rank (0.75) versus 0.60 for quantile and 0.48 for equal-width. In the majority of datasets, the two best discretizers (usually quantile and k-means) are statistically tied. Neither quantile nor uniform outperform k-means at the 5% level on any dataset. However, k-means wins outright in 3 datasets, most strikingly by 55%, on the Brazilian Houses dataset. Its other wins range from 1-2%. Brazilian Houses is especially unique, because it contains the most highly skewed input dataset (averaged by column) among those in the benchmark. In addition, the extreme predictor values (e.g. high property tax) coincide with extreme target values (rent price), so isolating these points is crucial. In section 3.2, we conducted in-depth analysis into this relationship between k-means performance and highly-skewed data.
Zhang et al. (2017) recommend using a budget of when training GBDTs on a GPU; in Appendix 4, we rerun the regression suite to analyze this scenario. Under this tighter budget, k-means now holds an 8% lead on CPU Act and widens its lead on Brazilian Houses to 66%, while quantile gains a statistically significant 2.5% lead on Superconduct and a 0.5% lead on Diamonds.
@@
We also pre-binned three representative regression datasets with either quantile or our k-means cuts (B = 255) and trained XGBoost using its exact tree builder (no internal histogram). As Appendix 5 shows, k-means retains existing ties while reducing MSE on Brazilian Houses by 64%, confirming the lift translates unchanged to a mainstream GBDT engine when using exact no-histogram methods.
@@
| Dataset Name | Quantile | Uniform | k-means | Exact |

### v2

@@ Abstract
Modern Gradient Boosted Decision Trees (GBDTs) accelerate split finding with histogram-based binning, which reduces complexity from to by aggregating gradients into fixed-size bins. However, the predominant quantile binning strategy—designed to distribute data points evenly among bins—may overlook critical boundary values that could enhance predictive performance. In this work, we consider a novel approach that replaces quantile binning with a -means discretizer initialized with quantile bins, and justify the swap with a proof showing how, for any -Lipschitz function, k-means maximizes the worst-case explained variance of Y obtained when treating all values in a given bin as equivalent. We test this swap against quantile and uniform binning on 33 OpenML datasets plus synthetics that control for modality, skew, and bin budget. Across 18 regression datasets, k-means shows no statistically significant losses at the 5% level and wins in three cases—most strikingly a 55% MSE drop on one particularly skewed dataset—even though k-means’ mean reciprocal rank (MRR) is slightly lower (0.65 vs 0.72). On the 15 classification datasets the two methods are statistically tied (MRR 0.70 vs 0.68) with gaps 0.2 pp. Synthetic experiments confirm consistently large MSE gains—typically 20% and rising to 90% as outlier magnitude increases or bin budget drops. We find that k-means keeps error on par with exhaustive (no-binning) splitting when extra cuts add little value, yet still recovers key split points that quantile overlooks. As such, we advocate for a built-in bin_method=k-means flag, especially in regression tasks and in tight-budget settings such as the 32–64-bin GPU regime—because it is a "safe default" with large upside, yet adds only a one-off, cacheable overhead ( 3.5s per feature to bin 10M rows on one Apple M1 thread).
@@ 1 Introduction
Gradient Boosted Decision Trees (GBDTs) are ensemble learning methods that have achieved state-of-the-art results on a wide range of tasks. By iteratively fitting new decision trees to the "pseudo-residuals" of a loss function, GBDTs combine many weak learners (e.g. individual decision trees) into a strong predictor (Friedman, 2001). Since the seminal work of Friedman on gradient boosting, the technique has become a de facto choice for structured data problems. Because GBDTs require only a twice-differentiable loss function, they are especially useful for non-standard prediction problems, such as ranking or quantile regression (Burges, 2010). Modern implementations like XGBoost, LightGBM, and CatBoost have popularized GBDTs by offering order-of-magnitude computational speed-ups with excellent accuracy (Chen & Guestrin, 2016) (Ke et al., 2017) (Prokhorenkova et al., 2019). These systems have been widely adopted to win Kaggle and other ML competitions, often outperforming deep learning on tabular data (Grinsztajn et al., 2022).
In formal terms, GBDT algorithms wish to minimize an arbitrary twice-differentiable loss function by maintaining an ensemble-based predictor comprised of a series of CART trees. At each round t it calculates first and second-order statistics
and then fits a simple CART tree to maximize a split-gain criterion. The optimal weight for leaf is calculated as
and the model is updated as . Parameters and correspond to the learning rate and L2 regularization respectively, and are chosen by the user prior to training.
@@
A key method for optimizing these algorithms is their use of histogram-based binning for splitting continuous features. Traditional decision tree algorithms sort feature values to find split points, which is computationally expensive ( per feature initially, and for split-finding) and thus intractable for large datasets. To allow training on datasets sometimes too large to fit into memory, modern GBDT algorithms bucket continuous feature values into discrete bins and accumulate gradients per bin during training. This approximation drastically reduces computation and memory overhead while maintaining similar accuracy as exhaustive splitting. In particular, histogram-based models require only scans for histogram building (and only bins for split-finding), with no sorting needed. Major state-of-the-art algorithms, like XGBoost, LightGBM, and CatBoost, all use similar histogram-binning methods, though with small differences in implementation. These innovations allow GBDTs to effectively scale to datasets with millions or billions of instances.
@@
In state-of-the-art histogram-based GBDT libraries, the default choice is quantile binning: thresholds are placed so that every bin holds roughly the same number of observations, barring duplicate values. This equal-frequency distribution is popular because it preserves the rank order of values and often yields split quality comparable to those obtained by exhaustive-search algorithms. However, quantile binning can sometimes mask rare but influential values. For example, in a sample of one million observations with one thousand extreme outliers, a 255-bin quantile scheme would pool those outliers with 3000 ordinary points, blunting a potentially important boundary.
@@
A natural alternative is to utilize some method of unsupervised clustering to isolate important points prior to training. One such example is via 1-D k-means clustering (Lloyd, 1982), which explicitly minimizes within-bin variance and can place cuts at rare yet influential values. We therefore compare Lloyd–k-means clustering with quantile initialization111Appendix D benchmarks two other globally-optimal discretizers, MILP-optimal (Navas-Palencia, 2022) and 1-D -means (Wang & Song, 2011). These methods either have unreasonable computational requirements or merge features down to bins, significantly harming accuracy. with quantile binning as well as uniform (equal-width) binning to characterize the regimes—tail mass, multi-modality, and bin budget—under which each discretizer is preferable and to test whether a simple k-means swap can boost accuracy without compromising training speed.
@@
We find that uniform equal-width cuts seldom rival quantile, whereas k-means almost always draws level with quantile and occasionally provides a marked boost. Across 18 real-world regression datasets, k-means never performs worse than quantile at the 5% level and cuts MSE by more than half on the most skewed case, though it has a slightly lower MRR (0.65 vs 0.72). When we repeat the sweep with the 63-bin budget recommended for GPU training (Zhang et al., 2017), k-means posts five wins and raises its best-case drop to 68% MSE (Table 5), though quantile achieves two small () wins.
In 15 classification datasets quantile and k-means are statistically tied (MRR 0.68 vs 0.7) with gaps percentage points (pp)–differences that are essentially negligible. Synthetic regression benchmarks tell the same story: when outliers dominate or when bin budget is low, k-means often cuts error by more than 20% and can reduce error by as much as 90% in some cases. Taken together, our findings show that k-means stays within statistical noise of exhaustive splitting when the bin budget is generous, yet often uncovers critical split points that quantile misses.
To our knowledge, no open-source GBDT package (XGBoost, LightGBM, CatBoost) offers a k-means discretizer, and the literature lacks any systematic test of alternative unsupervised binning schemes. We therefore conduct the first large-scale benchmark (33 OpenML datasets plus controlled synthetics) and show that this drop-in option can significantly cut regression MSE in skewed datasets or in low-bin scenarios while adding only a one-off, cacheable preprocessing cost of 3.5s per feature to bin 10 M rows on a single Apple M1 thread—making a bin_method=k-means flag an immediately actionable upgrade for commercial GBDT libraries.
@@ 2 Related Work
In recent years, gradient-boosted decision trees (GBDTs) have become the dominant choice for large-scale tabular prediction thanks to highly optimized "histogram" implementations that reduce time complexity from to per feature by pre-binning continuous inputs. LightGBM introduced leaf-wise growth together with GPU-friendly, histogram-based training to parallelize histogram-building and split-finding on large datasets (Ke et al., 2017). XGBoost introduces a weighted quantile sketch to obtain approximate split points that are -accurate even on distributed data (Chen & Guestrin, 2016). CatBoost further refines this pipeline with ordered boosting and efficient handling of categorical features, yielding state-of-the-art accuracy on many public benchmarks (Prokhorenkova et al., 2019). Scikit-learn, a package for easy application of ML algorithms, recently added HistGradientBoosting as a re-implementation of these histogram algorithms for GBDTs, providing a convenient baseline for academic comparisons (Pedregosa et al., 2018).
Although the histogram abstraction is now standard, the binning step itself has attracted surprisingly little study. All three major libraries—XGBoost, LightGBM, and CatBoost—still build their histograms from equal-frequency (quantile) cuts. Their main differences lie elsewhere: (1) XGBoost replaces the naïve per-feature sort with a weighted quantile sketch—a streaming algorithm that yields -accurate percentiles in time and sub-linear memory, allowing for distributed quantile binning (Chen & Guestrin, 2016); (2) LightGBM applies gradient-based one-side sampling (GOSS) after bins are formed, keeping rows with large gradients to speed up tree growth (Ke et al., 2017); (3) CatBoost introduces ordered boosting and specialized handling of categorical features (Prokhorenkova et al., 2019). Thus, despite diverse engineering optimizations, the underlying assumption that equal-frequency bins are “good enough” remains largely unchallenged.
Adaptive or data-driven binning has been explored in other settings. Early entropy-based cuts (Fayyad & Irani, 1993) and MDL-guided binning (Dougherty et al., 1995) found that supervised partitions improve decision-tree and Naïve-Bayes accuracy over equal-width binning in classification settings. In federated settings, Ong et al. (2020) begin with equal-mass bins and dynamically merge or split them using gradient entropy to keep communication costs fixed. Our work is largely orthogonal to these settings: we study library-level 222By “library-level” we refer to modifications that live inside the GBDT library itself (e.g. LightGBM, XGBoost). The only exposure to the end user is an extra option to specify the bin_method – no separate preprocessing script or data-pipeline changes are required., unsupervised binning that can serve as the starting point for any histogram-GBDT pipeline, federated or centralized333However, we do conduct a small benchmarking experiment in Appendix D which includes a supervised binning method.. Although k-means is unsupervised, its variance-minimizing objective lets it capture label-relevant tails that equal-mass schemes often smooth over, thus recovering a portion of the gains previously attributed only to supervised or adaptive methods.
@@
3 Theoretical Motivation
Although our contribution is primarily empirical, we provide a short analysis showing that for any -Lipschitz target , -means binning maximizes the tightest possible lower bound on the explained variance of obtained when treating all values in a given bin as equivalent.
Throughout this section we work in the population (distributional) setting: a -binning is a measurable partition of the feature domain, and we write . Given centers the -means quantization (binning) objective is to minimize
For any fixed partition , the objective is minimized with respect to the centers when ; substituting yields
Because the optimization is taken jointly over all measurable and the centers , population-level -means is therefore exactly the problem of minimizing the expected within-bin variance across all partitions. This mirrors the familiar sample version, where -means minimizes the within-cluster variance of the observed data points (Lloyd, 1982).
Lemma 1 (Paired–difference identity).
Let be a square-integrable random variable and let be an independent copy of . Then
Proof.
Theorem 1.
Let random variables (X, Y) with finite second moments satisfy a -Lipschitz continuous model with . Then, k-means binning chooses the bins which maximize the tightest possible lower bound on the explained variance of Y.
Proof.
Fixing integer K and measurable binning B as , define . We write and as the expectation/variance of some function of g across bins under .
For regression, we have that the explained variance of a given binning B is
By the law of total variance. Since Y and X are fixed before bin choice, the explained variance is maximized when is minimized. Using Lemma 1 and the -Lipschitz property of f, we have that
Consequently,
Since -means minimizes (as shown earlier), it minimizes this upper bound and therefore maximizes a lower bound on the explained variance of Y. Since this bound is strict when (proof in Appendix A), this is the tightest possible lower bound on the explained variance of Y.
Since exact k-means binning is impractical for large datasets444Appendix D presents a brief benchmark of the 1-D dynamic-programming formulation of optimal -means (Wang & Song, 2011). The method’s computational cost is so high that it becomes impractical for histogram-based GBDTs, offering no realistic speed–accuracy advantage over Lloyd’s algorithm., in Section 4 we replace it with Lloyd’s algorithm via scikit-learn, which offers a fast heuristic approximation to the exact . We find that this theoretical justification carries over when using Lloyd’s algorithm on real-world datasets, with k-means binning often obtaining the lowest realized error.
@@ 4 Experiments
We evaluate our discretizers on 33 OpenML tasks and a suite of controlled synthetic benchmarks. An anonymized reproducibility package—source code, logs, and result tables—is provided for reviewers (see the Links section).
@@ 4.1.1 Methodology
To ensure replicable results we evaluate our models on the OpenML (Vanschoren et al., 2014) benchmark suite described in Grinsztajn et al. (2022) (study_id 336 for regression, 337 for binary classification), dropping one task from each track (Higgs, Zurich Delays) due to computational constraints. The remaining 18 regression and 15 classification tasks span – instances and 2–420 numeric features. Appendix B lists observations and features for each dataset.
@@
We compare three discretizers: quantile (LightGBM default), uniform (equal-width), -means (Lloyd with quantile seeding). Unless noted otherwise, all binning schemes use bins (the common default in LightGBM and scikit-learn, among others).555Appendix E reports an ablation at the GPU-recommended budget of , as described in Zhang et al. (2017).
In all real-world benchmark tables, we also list an "exhaustive search" column, which checks all possible split thresholds without using binning. Because that variant enumerates all possible split thresholds (O(N) per node), it is impractically slow on large datasets but serves as a practical upper-bound reference to quantify the accuracy degradation attributable to binning itself.
In Appendix D, we present a brief benchmark of two further binning schemes and show that neither achieves an attractive speed–accuracy trade-off.
@@
Commercial-grade libraries such as LightGBM, XGBoost, and CatBoost apply additional preprocessing layers—e.g. LightGBM’s exclusive-feature bundling, or XGBoost’s weighted quantile sketch (Ke et al., 2017) (Chen & Guestrin, 2016). These steps are re-executed even when the input has already been pre-binned to 255 distinct values, producing a second, library-specific histogram that blurs the effect we wish to measure. To isolate the contribution of the external discretizer itself, we therefore run all main-paper experiments with scikit-learn’s vanilla GradientBoostingRegressor/Classifier, where the model consumes our bins exactly as supplied. In a supplementary run with XGBoost (Appendix F)—configured to bypass its internal quantile sketch—we observed similar results on representative regression datasets, suggesting that the choice of baseline model does not influence our conclusions.
For every (dataset, binning, learner) triple we run a 30-trial RandomizedSearchCV with 5-fold CV on hyper-parameters shown in Appendix C. Each experiment is repeated over 20 random train/test splits (80/20) to estimate variability.
@@
We report mean squared error (MSE) for regression and ROC-AUC for binary classification. Regression MSE values span several orders of magnitude, so each row of our results table presents them in scientific notation with the common exponent factored out on the left of the dataset name (e.g. "Brazilian Houses "). This keeps the numeric columns directly comparable across datasets.
To provide a scale-free summary metric, we compute a macro-averaged mean-reciprocal-rank (MRR, is better). For each dataset the three histogram methods (exhaustive search excluded) are ranked; the reciprocal rank is then averaged over datasets, with each dataset contributing one vote.
Paired two-sided t-tests (n=20 splits) compare the top-ranked discretizer with the runner-up (exhaustive search excluded) on each dataset. All resulting p-values are jointly adjusted with the Benjamini–Hochberg procedure (Benjamini & Hochberg, 1995); cells that stay significant at the level after this correction are bolded.
@@ 4.1.2 Results
Tables LABEL:tab:regression_datasets and LABEL:tab:classification_datasets report mean test-set scores averaged over 20 random 80:20 splits, with per-dataset winners flagged by paired t-tests, as described in section 4.1.1.
Regression (18 tasks) Although k-means attains a slightly lower MRR compared to quantile binning (0.65 vs. 0.72), it is never statistically outperformed by either quantile or uniform binning at the 5% significance level across all evaluated datasets. However, k-means wins outright in 3 datasets, most strikingly by 55%, on the Brazilian Houses dataset. Its other wins range from 1-2%. Brazilian Houses is especially unique, because it contains the most highly skewed input dataset (averaged by column) among those in the benchmark. In addition, the extreme predictor values (e.g. high property tax) coincide with extreme target values (rent price), so isolating these points is crucial. Across the full regression suite, -means performs worse than the exhaustive split method on just three datasets; of those, quantile performs statistically significantly worse than -means on two (and significantly better on none). This suggests that -means retains almost all of the predictive value of exhaustive splitting when extra cuts add little value, yet still recovers key tail-driven split-points that quantile overlooks. In section 4.2, we conducted in-depth analysis into this relationship between k-means performance and highly-skewed data.
Zhang et al. (2017) recommend using a budget of when training GBDTs on a GPU; in Appendix E, we rerun the regression suite to analyze this scenario. Under this tighter budget, k-means now holds an 8% lead on CPU Act and widens its lead on Brazilian Houses to 66%, while quantile gains a statistically significant 2.5% lead on Superconduct and a 0.5% lead on Diamonds. In Appendix E, we hypothesize two mechanisms for quantile’s small wins in the low-bin regime: (1) k-means may spend a larger proportion of a small bin budget isolating extremes that may not be label-relevant, and (2) coarser histograms also introduce greater variance, so edge-label alignment can occur by chance.
@@
We also pre-binned three representative regression datasets with either quantile or our k-means cuts (B = 255) and trained XGBoost using its exact (i.e. exhaustive) tree builder. As Appendix F shows, k-means retains existing ties while reducing MSE on Brazilian Houses by 64%, confirming the lift translates unchanged to a mainstream GBDT engine when using exhaustive no-histogram methods.
@@
| Dataset Name | Quantile | Uniform | k-means | Exhaustive |